LHC 2010 - latest news

Week 20: Machine Coordinators:    Oliver Bruening - Joerg Wenninger

Latest plan: link PDF link XLS (version of 14 May, 13:30)

List of End of Fill studies (in work)

List of MD requests (in work)

Procedure for Physics fill and recover including squeeze

Day Short summary/plan Link to reports from 8:30 and 17:00 meetings
Monday 24th May
  • 00h30: Beam lost - LBDS dump.
  • 03h00: Injection. Again LBDS dump.
  • 04h00: Access for LBDS
  • 09h00: Injection.
Report from 9h00 meeting
Sunday 23rd May
  • 06h00: Stable beams
  • 12h30: Beam lost - LBDS, tracking error.
  • 16h00: Injecting.
  • 21h30: Stable beams
Report from 9h00 meeting
Saturday 22nd May
  • 04h00: Stable beams
  • 11h30: Beam lost PC trip.
  • 13h00: Access for PC repair.
  • 20h00: Precycle.
  • 21h40: Preparing for physics with 6x2E10, b* = 2m
  • 23h30: Ramp. Lost beams on flat top.
Report from 9h00 meeting
Friday 21st May
  • 08h00: RF capture studies at injection.
  • 12h00: LBDS studies.
  • 19:00: Preparing for physics with 6x2E10, b* = 2m
  • 22:30: beam lost in ramp. RT trims lead to 25 COD trips.
Report from 8h30 meeting
Thursday 20th May
  • 02h00: Aperture and off-momentum beta-beat at injection.
  • 06h00: Cryo de-icing S45.
  • 08h00: Access....
  • 18h00-23h00: QPS tests S67 10 A/s.
Report from 8h30 meeting
Wednesday 19th May 2010
  • 03h00: Prepare for stable beams with 6x2e10 per beam.
  • 07h30: Beams lost after trip of RCBX3.L1.
  • 09h00: Access.
  • 10h00: Recover...
  • 13h00: injection and ramp 1e11 for beam stability studies. Beam lost in the ramp.
  • 19h00: Damper and orbit FB studies at injection, test ramp.
Report from 8h30 meeting
Tuesday 18th May 2010
  • 00h00: Recovery, prepare for stable beams with 6x2e10 per beam.
  • 05h00: stable beams.
  • 05h40: Beam dumped, electrical network perturbation.
  • 09:00: Access.
  • 12:00: injection and ramp 1e11 for collimation setup.
  • 17:40: Beams lost.
  • 20:00 : Prepare physics fill.
  • 23:30 : lost during squeeze due to COD trips.
Report from 8h30 meeting
Monday 17th May 2010
  • 00h07: Recovery and injection setup: 6x2e10 per beam.
  • 06h00: Stable beams (6 x 2e10 / beam)
  • 14h30: dump and recover
  • 18h00: injection and ramp 1e11 to study the (in)stability.
Report from 8h30 meeting